# Dislocation correlations in GaN epitaxial films revealed by EBSD and XRD

### v1

@@ 4 Discussion
4.1 Strain distributions in XRD and EBSD
Both XRD profiles and probability distributions obtained from EBSD maps represent distributions of strains and rotations, as it was already pointed out and compared by Wilkinson et al. [15]. Measurements by these two methods are complimentary. The XRD intensity provides the probability distribution over several orders of magnitude down from the maximum, allowing the dislocation density $\varrho$ and the dislocation strain screening distance $R$ to be determined from a single profile. However, the XRD intensity is an average over the sample volume and potentially includes a contribution from strained regions far from the surface, in particular the film–substrate interface. EBSD, in turn, provides the maps of strains and rotations at the surface. The statistics of the strain maps are limited by the number of pixels in the map. The maps can be used to determine not only the strain probability distributions, but also the spatial correlations of strains and rotations, which allows the screening distance $R$ to be determined directly as the distance at which the correlations become zero. A comparison of the dislocation correlations obtained by the two methods is the main aim of the present paper.
We have compared in Fig. 2, the measured XRD curves, their Monte Carlo simulations, and fits of both sets of curves by the formula proposed in Ref. [30]. A comparison of a curve simulated by the Monte Carlo method with the fit of this curve provides a check on the internal consistency of simulations and fits: dislocation densities obtained from fits of simulated curves (open symbols in Fig. 2(c)) coincide with the input values of the simulations with fairly little scatter.
The dislocation densities that are obtained in the Monte Carlo simulations and confirmed by the fits of the XRD curves are larger than the spot densities obtained from Fig. 1 by factors of 1.6 and 2.7 for samples 1 and 2, respectively. It is noteworthy that our previous studies [30, 37] also yielded dislocation densities obtained from the XRD profiles that were several times larger than those determined by transmission electron microscopy. Since XRD collects intensity scattered from the whole volume of the GaN film, the additional broadening of XRD curves can be caused by strain originating from the film–substrate interface. However, calculations [37] show that misfit dislocations at the film–substrate interface modify the reciprocal space maps but, for thick films, have little effect on the double-crystal curves that are measured in the present work.
@@ 4.2 Screening distances
The screening of dislocation strains by surrounding dislocations provides a transition from the Gaussian central part of the XRD profile to the $I(q)\propto q^{-3}$ tails. The screening distance $R$ is a second parameter, in addition to the dislocation density $\varrho$, that is included in the fit of the XRD profiles [30]. The screening distances obtained by fitting the experimental curves (full symbols in Fig. 2(d)) agree with those obtained from correlations in EBSD maps (horizontal dashed lines in Fig. 2(d)), although there is a rather large scatter between the values obtained from different reflections.
Plots of the autocorrelation functions of the components of the strain and rotation tensors measured by EBSD provide a more robust tool for determining the screening distance $R$. This distance is clearly seen as a kink in the plot of the correlation function in dependence of the logarithm of the separation, and is found to be the same for all strain and rotation components. We find screening distances $R$ of 2 µm and 0.3 µm for the GaN films with threading dislocation densities $\varrho$ of $5\times 10^{8}$ cm-2 and $1.8\times 10^{10}$ cm-2, respectively. In both cases, the number of dislocations that provide screening of the strain of a given dislocation $M=R\varrho^{1/2}$ is about 4. This means that the dislocation strain field is very effectively screened by only a few surrounding dislocations. To obtain these screening distances, we measure the strain and rotation maps over areas with linear sizes that are large compared to $R$ using steps that are small compared to $R$. Different map sizes and steps are chosen depending on the dislocation density. We note that very close numbers, 4 to 6 dislocations providing screening of the dislocation strain fields, are obtained for plastically deformed metals [38] by an XRD line profile analysis method that is close to our approach for GaN films.
@@ 4.3 Geometrically necessary dislocations
Maps of the strain and rotation components are commonly used in EBSD studies to determine the densities of geometrically necessary dislocations (GND) [39, 40, 41, 42, 43, 44]. The GND density corresponds to the minimum density of dislocations that is required to provide a net effect that is of interest [45, 46]. It is therefore dependent on the scale of that effect and implies an appropriate spatial averaging of the strains and rotations at smaller scales. The papers cited above use the concept of GND to describe boundaries between misoriented domains or plastic bending of a crystal. The GND value is then a minimum density of dislocations that provides a small angle boundary or the curvature of the crystal lattice. In these cases, strain variations can be neglected compared to the lattice rotations, and the GND can be obtained from the differences in rotations between neighboring points in the measured maps.
@@
4.4 Spatial resolution of EBSD maps
@@
We find that the correlation functions deviate from the $\propto\ln r$ law at small distances and become more gradual. Also, the EBSD maps of sample 2 with high dislocation density measured with a step between the measuring points of only 20 nm, show a clear distinction between the $x$ and $y$ directions. A comparison of the correlations in the $x$ and $y$ directions and Monte Carlo simulations allow us to estimate the spatial resolution of the EBSD maps to be $20\times 200$ nm2, with the worse resolution along the direction of the tilt of the incident electron beam. In metals, grain boundaries are used for a direct measurement of the resolution, which depends on the material and the accelerating voltage (we refer to the papers [47, 48] which also provide literature reviews). The resolutions in $x$ and $y$ directions are obtained by appropriately orienting the grain boundary with respect to the direction of the electron beam. The resolution that we obtain is comparable to that obtained in relatively light metals but shows a larger anisotropy between the $x$ and $y$ directions.
@@
For additional insight, we simulate the trajectories of the electrons in GaN using the free software CASINO [49]. Figure 7(a) is a side view on the electron trajectories. Red lines are trajectories of the backscattered electrons. A top view of their exit points is shown in Fig. 7(b). Their spatial distribution is very asymmetric with respect to the $x$ and $y$ directions. The projections of the outcrop distributions on the $x$ and $y$ axes in Figs. 7(c,d) are even broader than the resolution that we obtain from the simulations of the strain and rotation maps. Since low-loss electrons provide the major contribution to EBSD patterns [50], we plot by dotted lines in Figs. 7(c,d) the distributions of electrons that lost less than 10% of their initial energy of 15 keV. The widths of these distributions are consistent with the resolution that we found.
@@
The EBSD maps of the dislocation-free GaN sample 0 show only uncorrelated random noise for all strain and rotation components. The widths of these distributions are an order of magnitude larger compared to Si(001). We also note that the XRD curves of this sample, albeit narrow compared to the other samples, are significantly broader than expected for its dislocation density [19]. The EBSD measurements on GaN sample 0 were performed before and after surface cleaning, which had a negligible effect on the strain distributions. Hence, neither surface roughness nor contamination do appear to be responsible for the apparent strain in this sample.
@@
In the absence of dislocations, the remaining source of strain are point defects, including impurities as well as native defects. The density of native defects in GaN is orders of magnitude larger than in Si. The density of Ga vacancies alone in GaN ranges from $10^{16}$$10^{18}$ cm-3 [51, 52, 53], as compared to a vacancy density below $1\times 10^{14}$ cm-3 in Si [54, 55]. To examine the effect of point defects, we have performed simulations of their strain maps, assuming a spatially random distribution of dilatation centers with the excess volume equal to the atomic volume. Even for a density of $10^{18}$ cm-3, the strain produced by these centers is too low to account for the experimentally measured strain in Fig. 3(a). Moreover, the simulated strain maps exhibit the same anisotropy as seen in Fig. 3(c) caused by the inclination of the incident electron beam.
@@ 5 Summary
The screening of the dislocation strains by the surrounding dislocations is revealed by both fits of the XRD curves in reciprocal space and the real space correlations in the EBSD strain maps. It is shown that the characteristic screening distance can be derived with a higher accuracy from the EBSD strain and rotation maps than from XRD profiles. Screening distances of 2 µm and 0.3 µm are obtained for the GaN films with threading dislocation densities of $5\times 10^{8}$ cm-2 and $1.8\times 10^{10}$ cm-2, respectively, which indicates that the dislocation strain is screened by only 4 neighboring dislocations in both samples. The strain maps from the high dislocation density sample measured with a small step size show a large anisotropy in the resolution of the EBSD measurements. The resolution is estimated to be $20\times 200$ nm2, with the worse resolution in the direction corresponding to the inclination of the incident electron beam. XRD curves and EBSD maps are simulated by the Monte Carlo method within one and the same model of dislocation distributions. We find that XRD overestimates the dislocation densities, presumably due to contributions from the strained region near the film–substrate interface. In contrast, EBSD measurements on a sample with a high dislocation density underestimate the strain, presumably due to large strain gradients within the probed volume.
@@ Acknowledgments
The authors thank Bernd Jenichen for XRD measurements on sample 2 made for our former study [19], Jingxuan Kang for cleaning sample 0 and Moritz Hansemann for AFM measurements of that sample, as well as Achim Trampert for a critical reading of the manuscript.
Supplementary Information
Figure S1: Kikuchi patterns of the GaN samples 0, 1, 2 and the Si wafer.
Figure S2: Components of the strain and rotation tensors of sample 0 (dislocation free GaN sample).
Figure S3: Components of the strain and rotation tensors of sample 1.
Figure S4: Components of the strain and rotation tensors of sample 2.
Figure S5: Components of the strain and rotation tensors of a silicon wafer. Note different ranges of the values in comparison to Fig. S2.
Figure S6: Monte Carlo simulation of the strain and rotation tensors of sample 1.
Figure S7: Monte Carlo simulation of the strain and rotation tensors of sample 2.
Figure S8: Orientation-averaged autocorrelation functions of the rotations (a,b) in the measured maps and (c,d) in Monte Carlo simulations.

### v2

@@ 4 Discussion
4.1 Strain distributions in XRD and HR-EBSD
Both XRD profiles and probability distributions obtained from HR-EBSD maps represent distributions of strains and rotations, as it was already pointed out and compared by Wilkinson et al. [27]. Measurements by these two methods are complimentary. The XRD intensity provides the probability distribution over several orders of magnitude down from the maximum, allowing the dislocation density $\varrho$ and the dislocation strain screening distance $R$ to be determined from a single profile. However, the XRD intensity is an average over the sample volume and potentially includes a contribution from strained regions far from the surface, in particular the film–substrate interface. HR-EBSD, in turn, provides the maps of strains and rotations at the surface. The statistics of the strain maps are limited by the number of pixels in the map. The maps can be used to determine not only the strain probability distributions, but also the spatial correlations of strains and rotations, which allows the screening distance $R$ to be determined directly as the distance at which the correlations become zero. A comparison of the dislocation correlations obtained by the two methods is the main aim of the present paper.
We have compared in Fig. 2 the measured XRD curves, their Monte Carlo simulations, and fits of both sets of curves by the formula proposed in Ref. [42]. A comparison of a curve simulated by the Monte Carlo method with the fit of this curve provides a check on the internal consistency of simulations and fits: dislocation densities obtained from fits of simulated curves (open symbols in Fig. 2(c)) coincide with the input values of the simulations with fairly little scatter.
The dislocation densities that are obtained in the Monte Carlo simulations and confirmed by the fits of the XRD curves are larger than the spot densities obtained from Fig. 1 by factors of 2 and 3 for samples 1 and 2, respectively. It is noteworthy that our previous studies [42, 54, 55] also yielded dislocation densities obtained from the XRD profiles that were several times larger than those determined by transmission electron microscopy. Since XRD collects intensity scattered from the whole volume of the GaN film, the additional broadening of XRD curves can be caused by strain originating from the film–substrate interface. However, calculations [55] show that misfit dislocations at the film–substrate interface modify the reciprocal space maps but, for thick films, have little effect on the double-crystal curves that are measured in the present work.
@@ 4.2 Screening distances
The screening of dislocation strains by surrounding dislocations provides a transition from the Gaussian central part of the XRD profile to the $I(q)\propto q^{-3}$ tails. The screening distance $R$ is a second parameter, in addition to the dislocation density $\varrho$, that is included in the fit of the XRD profiles [42]. The screening distances obtained by fitting the experimental curves (full symbols in Fig. 2(d)) agree with those obtained from correlations in HR-EBSD maps (horizontal dashed lines in Fig. 2(d)) within a factor of 2, with a rather large scatter between the values obtained from different reflections.
Plots of the autocorrelation functions of the components of the strain and rotation tensors measured by HR-EBSD provide a more robust tool for determining the screening distance $R$. This distance is clearly seen as a kink in the plot of the correlation function in dependence of the logarithm of the separation, and is found to be the same for all strain and rotation components. We find screening distances $R$ of 2 µm and 0.3 µm for the GaN films with threading dislocation densities $\varrho$ of $5\times 10^{8}$ cm-2 and $1.8\times 10^{10}$ cm-2, respectively. In both cases, the number of dislocations that provide screening of the strain of a given dislocation $M=R\varrho^{1/2}$ is about 4. This means that the dislocation strain field is very effectively screened by only a few surrounding dislocations. To obtain these screening distances, we measure the strain and rotation maps over areas with linear sizes that are large compared to $R$ using steps that are small compared to $R$. Different map sizes and steps are chosen depending on the dislocation density. We note that very close numbers, 4 to 6 dislocations providing screening of the dislocation strain fields, are obtained for plastically deformed metals [56] by an XRD line profile analysis method that is close to our approach for GaN films.
@@ 4.3 Geometrically necessary dislocations
Maps of the strain and rotation components are commonly used in HR-EBSD studies to determine the densities of geometrically necessary dislocations (GND) [57, 58, 59, 60, 61, 62]. The GND density corresponds to the minimum density of dislocations that is required to provide a net effect that is of interest [63, 64]. It is therefore dependent on the scale of that effect and implies an appropriate spatial averaging of the strains and rotations at smaller scales. The papers cited above use the concept of GND to describe boundaries between misoriented domains or plastic bending of a crystal. The GND value is then a minimum density of dislocations that provides a small angle boundary or the curvature of the crystal lattice. In these cases, strain variations can be neglected compared to the lattice rotations, and the GND can be obtained from the differences in rotations between neighboring points in the measured maps.
@@
4.4 Spatial resolution of HR-EBSD maps
@@
We find that the correlation functions deviate from the $\propto\ln r$ law at small distances and become more gradual. Also, the HR-EBSD maps of sample 2 with high dislocation density measured with a step between the measuring points of only 20 nm, show a clear distinction between the $x$ and $y$ directions. A comparison of the correlations in the $x$ and $y$ directions and Monte Carlo simulations allow us to estimate the spatial resolution of the HR-EBSD maps to be $20\times 200$ nm2, with the worse resolution along the direction of the tilt of the incident electron beam. In metals, grain boundaries are used for a direct measurement of the resolution, which depends on the material and the accelerating voltage (we refer to the papers [65, 66] which also provide literature reviews). The resolutions in $x$ and $y$ directions are obtained by appropriately orienting the grain boundary with respect to the direction of the electron beam. The resolution that we obtain is comparable to that obtained in relatively light metals but shows a larger anisotropy between the $x$ and $y$ directions.
@@
For additional insight, we simulate the trajectories of the electrons in GaN using the free software CASINO [67]. Figure 7(a) is a side view on the electron trajectories. Red lines are trajectories of the backscattered electrons. A top view of their exit points is shown in Fig. 7(b). Their spatial distribution is very asymmetric with respect to the $x$ and $y$ directions. The projections of the outcrop distributions on the $x$ and $y$ axes in Figs. 7(c,d) are even broader than the resolution that we obtain from the simulations of the strain and rotation maps. Since low-loss electrons provide the major contribution to HR-EBSD patterns [68], we plot by dotted lines in Figs. 7(c,d) the distributions of electrons that lost less than 10% of their initial energy of 15 keV. The widths of these distributions are consistent with the resolution that we found.
@@
The HR-EBSD maps of the dislocation-free GaN sample 0 show only uncorrelated random noise for all strain and rotation components. The widths of these distributions are an order of magnitude larger compared to Si(001). We also note that the XRD curves of this sample, albeit narrow compared to the other samples, are significantly broader than expected for its dislocation density [31]. The HR-EBSD measurements on GaN sample 0 were performed before and after surface cleaning, which had a negligible effect on the strain distributions. Hence, neither surface roughness nor contamination do appear to be responsible for the apparent strain in this sample.
@@
In the absence of dislocations, the remaining source of strain are point defects, including impurities as well as native defects. The density of native defects in GaN is orders of magnitude larger than in Si. The density of Ga vacancies alone in GaN ranges from $10^{16}$$10^{18}$ cm-3 [69, 70, 71], as compared to a vacancy density below $1\times 10^{14}$ cm-3 in Si [72, 73]. To examine the effect of point defects, we have performed simulations of their strain maps for a spatially random distribution of dilatation centers with the excess volume equal to the atomic volume. Even for a density of $10^{18}$ cm-3, the strain produced by these centers is too low to account for the experimentally measured strain in Fig. 3(a). Moreover, the simulated strain maps exhibit the same anisotropy as seen in Fig. 3(c) caused by the inclination of the incident electron beam.
@@ 5 Summary
The screening of the dislocation strains by the surrounding dislocations is revealed by both fits of the XRD curves in reciprocal space and the real space correlations in the HR-EBSD strain maps. Our study is performed using a well-defined array of threading dislocations in GaN(0001) as an example. It allowed Monte Carlo simulations of XRD curves and HR-EBSD maps within one and the same model of dislocation distributions and its quantitative comparison with the experiment. The simulations provide a basis for determining of the screening distances from the HR-EBSD maps as the positions of the kinks in the strain autocorrelations plotted against the logarithm of the distance. Determining the screening distance in this way does not require any simulations or fits and provides a robust tool also for more complicated dislocation ensembles, where simulations would be more difficult.
In the XRD, dislocation densities and screening distances are obtained by fitting the diffraction profiles. We find that XRD overestimates the dislocation densities, presumably due to contributions from the strained region near the film–substrate interface. The screening distances obtained by XRD from different reflections possess larger scattering and are consistent with those obtained by HR-EBSD within a factor of two.
Screening distances of 2 µm and 0.3 µm are obtained for the GaN films with threading dislocation densities of $5\times 10^{8}$ cm-2 and $1.8\times 10^{10}$ cm-2, respectively. This indicates that the dislocation strain is screened by only four neighboring dislocations in both samples. The strain maps from the high dislocation density sample measured with a small step size show a large anisotropy in the resolution of the HR-EBSD measurements. The resolution is estimated to be $20\times 200$ nm2, with the worse resolution in the direction corresponding to the inclination of the incident electron beam.
@@ Acknowledgments
The authors thank Bernd Jenichen for XRD measurements on sample 2 made for our former study [31], Carsten Richter for useful discussions, Jingxuan Kang for cleaning sample 0 and Moritz Hansemann for AFM measurements of that sample, as well as Achim Trampert for a critical reading of the manuscript. Arno Winkelmann and Graham Meaden read the preprint of the present paper and analyzed our HR-EBSD data with their software. We are grateful to them for their efforts and comments.
Supplementary material to the paper
Dislocation correlations in GaN epitaxial films
revealed by EBSD and XRD
by
Vladimir M. Kaganer, Domenik Spallek, Philipp John, Oliver Brandt, and Jonas Lähnemann
SM  1 Kikuchi patterns
Figure SM1: Kikuchi patterns of the GaN samples 0, 1, 2 and the Si wafer.
SM  2 Strain and rotation maps
Figure SM2: Components of the strain and rotation tensors of sample 0 (dislocation free GaN sample).
Figure SM3: Components of the strain and rotation tensors of sample 1.
Figure SM4: Components of the strain and rotation tensors of sample 2.
Figure SM5: Components of the strain and rotation tensors of a silicon wafer. Note different ranges of the values in comparison to Fig. SM2.
Figure SM6: Monte Carlo simulation of the strain and rotation tensors of sample 1.
Figure SM7: Monte Carlo simulation of the strain and rotation tensors of sample 2.
SM  3 Probability distributions and correlation functions of rotations
Figure SM8: Probability distributions of rotations (a,b) in the measured maps and (c,d) in Monte Carlo simulations.
Figure SM9: Orientation-averaged autocorrelation functions of the rotations (a,b) in the measured maps and (c,d) in Monte Carlo simulations.
SM  4 Effect of drift
Figure SM10 simulates the effect of drift on the correlation functions of strain. Figure SM10(a) is the Monte Carlo simulated map of the strain component $\varepsilon_{22}$ of sample 2. The same map is presented in Fig. 3(f) in the main text and in Fig. SM7. In Fig. SM10(b), a 15% drift in $y$-direction is simulated, assuming that the effect of drift is accumulated linearly: each pixel of the map is shifted vertically proportional to the time between the measurement of that pixel and the first pixel, so that the last pixel is shifted by 15%. Note that the drift effect is negligible at the top of the map and continuously increases to the bottom. The bottom part of the expanded map is cut out, to have the same area as in the initial map. Figure SM10(c) compares the autocorrelation functions obtained from these two maps. Since the correlations tend to zero at distances exceeding 0.2 µm, which is 0.1 of the map size, only correlations of strains in closely spaced points are of interest. As a result, the effect of drift on the autocorrelation function in Fig. SM10(c) is negligibly small.
Figure SM10: (a) Monte Carlo simulated map of the strain $\varepsilon_{22}$ of sample 2 (identical to Fig. 3(f) in the main text and Fig. SM7), (b) the same map with a 15% drift in $y$-direction added, and (c) autocorrelation functions obtained from these maps.
SM  5 Repeated measurements and Monte Carlo simulations of EBSD maps
Figure SM11: Autocorrelation functions of the strain component $\varepsilon_{22}$ obtained (a,b) in the HR-EBSD study of samples 1 and 2 and (c,d) by repeating Monte Carlo simulations of HR-EBSD maps with all parameters kept the same as in Figs. 3–6. Blue lines are the calculations in Fig. 6. Black lines in the experimental plots are due to the measurements in different places of the respective samples, while black lines in the Monte Carlo modeling are obtained by changing the seed of the random number generator. The vertical red arrows point out to the screening distances taken on input of the Monte Carlo simulations (2 µm and 0.3 µm for samples 1 and 2 respectively).